OpenStudy (anonymous):

Convert the following percent to a fraction62%

OpenStudy (ilovecomputers1):

62/100 which reduces to 31/50

OpenStudy (texaschic101):

this is how I do it... turn your percent to a decimal by moving decimal 2 spaces to the left. Giving us 0.62 now what place value is the last number 2 in ? It is in the 100dths place, so put it over 100, giving us 62/100 which reduces to 31/50. If your nuber would have been 0.235, then you would put 235 over 1000 because the last number, 5, is in the thousandths place. or just remember this...any percent is over 100.......20% = 20/100, 10% = 10/100, 50% = 50/100......and be sure to reduce your fraction.